OpenStudy (ttop0816):

Would my answer be correct?? There is a positive correlation between the number of times the Striped Ground Cricket chirps per second and the temperature in degrees Fahrenheit. If a scatter plot is made with the number of chirps on the horizontal axis and the trend line is found to be y = 3x + 25, then what would you predict the number of chirps per second to be when the temperature is 55 degrees Fahrenheit? 10<< i got this! 43 190

Miracrown (miracrown):

10, 42, 190 are the answer choices and you got 10?

OpenStudy (ttop0816):

@Miracrown yes!

Miracrown (miracrown):

How did you get 10? can't see that from the given info None of those choices actually make any sense based on the given info

OpenStudy (anonymous):

\[ y = 3x + 25\] \[55=3x+25\] looks good to me unless i am missing something

Miracrown (miracrown):

Oh, the answer is there but not 10

Miracrown (miracrown):

25, no?

OpenStudy (anonymous):

It is there and it seems good so far to me as well

OpenStudy (anonymous):

number of chirps is on the horizontal axis, it is the \(x\) value you are looking for

OpenStudy (ttop0816):

well i actually got two answers using two different ways! 10 & 190 How i got 10: y = 3x + 25 55 = 3x+25 30 = 3x How i got 190: y = 3x + 25 y = 3(55) + 25 y=190 I dont know which equation to use!!

OpenStudy (anonymous):

\[55=3x+25\\30=3x\\ 10=x\]

OpenStudy (ttop0816):

@satellite73 oh thank you!! (: i guess i did do it right~

OpenStudy (anonymous):

it is tricky so you have to read carefully

OpenStudy (anonymous):

The chirps per second is on the "horizontal" meaning the x-axis so can not be x

Miracrown (miracrown):

Oh, I see, it was my fault, I was reading it a different way

OpenStudy (anonymous):

it says the number of chirps is the "horizontal" axis

Miracrown (miracrown):

You're right...it's saying it's (temp) = 3(chirps) + 25 and then solve for chirps So y = 55 then x = 10

OpenStudy (anonymous):

The y-axis must represent the temperature

OpenStudy (ttop0816):

@Miracrown its alright(: but thanks for helping!

OpenStudy (anonymous):

so you are looking for \(x\)

Miracrown (miracrown):

yw @ttop0816